OpenStudy (dmezzullo):

Solve for x: 4 over 3x + 2 over 3 = 6 x = 4 x = 3 x = 1 x = 0

OpenStudy (dmezzullo):

x=3? @soty2013

OpenStudy (dmezzullo):

@sue101

OpenStudy (dmezzullo):

@jhonyy9

jhonyy9 (jhonyy9):

how is correct 4/3x +2/3 =6 hence is it right ?

jhonyy9 (jhonyy9):

the 3x is in denominator every two ?

OpenStudy (dmezzullo):

yes

jhonyy9 (jhonyy9):

ok than add to both sides -2/3 so than what will get ?

OpenStudy (dmezzullo):

4?

jhonyy9 (jhonyy9):

easy boy

jhonyy9 (jhonyy9):

so just step by step ok ?

OpenStudy (dmezzullo):

ok

OpenStudy (dmezzullo):

k

jhonyy9 (jhonyy9):

so than 4/3x +2/3 -2/3 =6 -2/3 ok ? can you continue it now ?

jhonyy9 (jhonyy9):

on the left part what will get ?

OpenStudy (dmezzullo):

2/3

jhonyy9 (jhonyy9):

2/3 -2/3 =0

jhonyy9 (jhonyy9):

so than will reamining there just 4/3x and on the right part ?

OpenStudy (dmezzullo):

ok

jhonyy9 (jhonyy9):

6-2/3 =(18-2)/3 =16/3 ok ?

jhonyy9 (jhonyy9):

so than there are 4/3x = 16/3 can you now cross multiplie and solve it for x

jhonyy9 (jhonyy9):

come on please because i need to go

OpenStudy (dmezzullo):

is the asnwer 4?

jhonyy9 (jhonyy9):

cross multiplie and will get 12=48x right ?

OpenStudy (dmezzullo):

yes

OpenStudy (dmezzullo):

and 48/12=4

jhonyy9 (jhonyy9):

no not is because you have wrote in the first lines that x is in denominator with 3

OpenStudy (dmezzullo):

what?

jhonyy9 (jhonyy9):

but seeing your choises you can getting it just when the x is in numerator

jhonyy9 (jhonyy9):

what mean that your exercise is 4x/3 +2/3 =6

OpenStudy (dmezzullo):

OMG IM SO FREAKIN LOST

jhonyy9 (jhonyy9):

so than 4x+2=18 4x=16 x=4 so this is the right solution for your exercise ok?

OpenStudy (dmezzullo):

k